OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + x2 – 20x ? 0, –5, –4 0, –5, 4 0, 5, –4 0, 5, 4

OpenStudy (anonymous):

@KirbyLegs could you help me out again please?

OpenStudy (anonymous):

If i get a medal . . .jk XD

OpenStudy (anonymous):

already gave you it lol

OpenStudy (anonymous):

okay lol

OpenStudy (anonymous):

HI :D

OpenStudy (anonymous):

ok could you explain how to find out the zeros cause idk how thats done

OpenStudy (anonymous):

The answers are 0,5, and -4. I'll show you why:

OpenStudy (anonymous):

hi @sourwing

OpenStudy (anonymous):

First you should simplify x(x^2 + x - 20)

OpenStudy (anonymous):

You can then do the opposite of FOIL x(x+5)(x-4) x=0 x+5=0 x=-5 x-4=0 x=4

OpenStudy (anonymous):

Oh i see thank you. Could you help with another question? :) I'll give you a medal lol